# Resonant Multi-Scalar Production in the Generic Complex Singlet Model in the Multi-TeV Region

### v1

@@ IV Benchmark Points
  • S1: We apply the current limits, described in Section III, on $\sin\theta_{1}$ to produce expected maximum rates for the HL-LHC. The most constraining limits on $\sin\theta_{1}$ from Fig. 1 are used. Results from both the hard-cut method for scalar searches with Higgs precision data constraints and $\chi^{2}_{\rm Tot}$ fits from Eq. (41) are considered separately. In practice, the strongest constraints on $\sin\theta_{1}$ are found by considering the smallest maximum value of $|\sin\theta_{1}|$ from the ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=0$ and ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=1$ cases.
@@
The scenario S1 is intended to find maximum di-scalar resonance rates for the HL-LHC. Hence, the $h_{2}$ mass range $m_{i}+m_{j}\leq m_{2}\leq 1$ TeV is used for S1, where $m_{i}$ and $m_{j}$ are the masses of the $h_{2}$ decay products. Scenarios S2-S4 are used to project maximum rates for future $pp$ machines, such as the 100 TeV FCC-hh. In these scenarios we consider a mass range of $m_{i}+m_{j}\leq m_{2}\leq 12$ TeV. For scenarios S2-S4, we normalize to the maximum rates to the SM Higgs production rate at the mass of $h_{2}$. Hence, they are very generic and can be applied to many future collider proposals and many single $h_{2}$ production modes.
@@
While a range of $m_{2}$ are considered, only certain values of the mass of $h_{3}$ are used. These masses are chosen to open new decay channels for $h_{3}$, and possibly new phenomenology for the resonance searches. Assuming $\theta_{2}$ is negligible but non-zero, $h_{3}$ will have partial widths into SM gauge bosons and fermions:
@@
Whether or not these are the dominant branching ratios depends on the Higgs trilinear coupling $h_{1}-h_{1}-h_{3}$ which in turn depends on the parameter point that maximizes the $h_{2}\rightarrow h_{1}h_{3}$ or $h_{2}\rightarrow h_{3}h_{3}$ production rates. As we show below, for all benchmark points the branching ratio of $h_{3}\rightarrow h_{1}h_{1}$ is at or near one when kinematically allowed.
We now discuss the results for maximizing $h_{i}h_{j}$ resonant production under the various scenarios. After these results, we will comment on the decays of $h_{3}$ and the possibility of have three or four SM-like Higgs boson signals.
@@ IV.1.1 Scenario S1: HL-LHC projections
Figure 4: Benchmarks for HL-LHC from scenario S1, showing (a,c) maximum production rates for $pp\rightarrow h_{2}\rightarrow h_{i}h_{j}$ with the gluon fusion and VBF modes added together at $\sqrt{S}=14$ TeV, and (b,d) the corresponding branching ratios to different di-scalar states. Results from using the bounds from $\Delta\chi^{2}_{Tot}$ fits in Eq. (41) are shown in (a,b) and from the hard cut method in (c,d). Blue is for $h_{2}\rightarrow h_{1}h_{1}$, red for $h_{2}\rightarrow h_{1}h_{3}$, and green for $h_{2}\rightarrow h_{3}h_{3}$. Solid is for $m_{3}=130$ GeV, dotted for $m_{3}=200$ GeV, and dashed for $m_{3}=270$ GeV.
In figure 4 we show the upper bounds on (a,c) the maximum production rate for $pp\rightarrow h_{2}\rightarrow h_{i}h_{j}$ and (b,d) the corresponding branching ratios of $h_{2}\rightarrow h_{i}h_{j}$, $i,j=1,3$, for scenario S1. Figure 4(a,b) use constraints from a $\chi^{2}_{Tot}$ fit and (c,d) from the hard cut method for scalar searches with precision Higgs constraints. The maximum production rates and branching ratios are largely independent of the choice of $m_{3}$ once kinematically allowed.
For $m_{2}\lesssim 600$ GeV, the maximum branching ratios for $h_{2}\rightarrow h_{1}h_{3}$ and $h_{2}\rightarrow h_{3}h_{3}$ are near one, with the final state $h_{3}h_{3}$ slightly larger. For $h_{2}\rightarrow h_{1}h_{1}$ the maximum branching ratio is between $0.6$ and $0.8$. As can be seen in Eqs. (II-19), the values of the $\lambda_{123}$ and $\lambda_{233}$ trilinear couplings have much more freedom from the choices of potential parameters than $\lambda_{112}$. Indeed, the ratio of the $h_{2}\rightarrow h_{1}h_{1}$ and $h_{2}\rightarrow ZZ$ partial widths are
@@
Hence, the partial width of $h_{2}\rightarrow h_{1}h_{1}$ cannot be too far from $h_{2}\rightarrow ZZ$ limiting how large the branching ratio of $h_{2}\rightarrow h_{1}h_{1}$ can be.
Once the mass of $h_{2}$ surpasses 600 GeV, all branching ratios decrease. The partial widths of $h_{2}$ into the massive SM gauge bosons $W^{\pm}W^{\mp}$ and $ZZ$ grow like the cubic power of $m_{2}$, whereas the partial widths into $h_{1}h_{3}$ and $h_{3}h_{3}$ do not. Additionally, the branching ratio of $h_{2}\rightarrow h_{1}h_{1}$ approaches the Goldstone boson equivalence theorem limit of ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})\approx 0.25$. Hence, the maximal branching ratio into $h_{i}h_{j}$ decrease in general. It is worth noting that the maximum branching ratio of the asymmetric decay $h_{2}\rightarrow h_{1}h_{3}$ decreases more quickly than the symmetric decay of $h_{2}\rightarrow h_{3}h_{3}$. In the small angle limit $|\theta_{1}|\ll 1$, the leading term of the trilinear $\lambda_{123}$ depends on trilinear couplings while $\lambda_{233}$ depends on quartic couplings. As we show in Appendix C, the combination of requiring the global minimum be the SM vacuum, perturbative unitarity, and a narrow width places an upper bound on $\lambda_{123}$ that is more stringent than the upper bound on $\lambda_{233}$. Indeed, as shown in Eq. (111), these constraints force the branching ratio of $h_{2}\rightarrow h_{1}h_{3}$ to approach zero as the mass of $h_{2}$ increases. While these derived constrains may not be the most stringent, they do help explain qualitatively why the branching ratio of $h_{2}\rightarrow h_{1}h_{3}$ decreases more quickly than the other di-scalar modes as $m_{2}$ increases.
@@ IV.1.2 Scenarios S2-S4: Future Collider Projections
The the fermionic partial widths only grow linearly with $m_{2}$ while the bosonic widths grow as $m_{2}^{3}$. Hence, $h_{2}\rightarrow f\bar{f}$, where $f$ are SM fermions, can be neglected and ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})\approx 0.25$.
@@
Two particularly surprising things occur when $m_{2}$ is in the multi-Tev range and above: the branching ratio of $h_{2}\rightarrow h_{3}h_{3}$ at its maximum rate approaches 0.5 and the maximum rates of $h_{2}\rightarrow h_{1}h_{1}$ and $h_{2}\rightarrow h_{3}h_{3}$ approach the same value. To understand ${\rm BR}(h_{2}\rightarrow h_{3}h_{3})\approx 0.5$, note that we are maximizing the rate:
@@
where we have use the narrow width approximation. Assuming the rate $pp\rightarrow h_{2}\rightarrow h_{3}h_{3}$ is maximized then ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})=0$ and the fermionic branching ratios are negligible. Equation (48) can be rewritten as
@@
A similar calculation can be made for $pp\rightarrow h_{1}h_{3}$, which explains why $|\sin\theta_{1}|$ at the maximum rate agrees for $h_{1}h_{3}$ and $h_{3}h_{3}$ production [see Fig. 3]. As discussed previously, theoretical constraints force ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})\rightarrow 0$ as $m_{2}$ becomes large. Hence, branching ratios and rates in Eqs. (52,53) cannot be saturated. There are constraints on $\lambda_{233}$ from global minimum, perturbative unitarity, and narrow width as shown in Eq. (101). However, these are significantly weaker than for $\lambda_{123}$, and the $pp\rightarrow h_{2}\rightarrow h_{3}h_{3}$ can saturate the bounds in Eqs. (52,53).
Figure 6: Benchmarks for future colliders, constrained by projections for HL-LHC and FCC-ee in scenario S3, showing (a) the maximum production times branching ratio scaled to the SM-like Higgs production rate and (b) the corresponding branching ratios. Blue is for $h_{2}\rightarrow h_{1}h_{1}$, red for $h_{2}\rightarrow h_{1}h_{3}$, and green for $h_{2}\rightarrow h_{3}h_{3}$. Solid is for $m_{3}=130$ GeV, dotted for $m_{3}=200$ GeV, and dashed for $m_{3}=270$ GeV
@@
The maximum rates for $pp\rightarrow h_{2}\rightarrow h_{3}h_{3}$ in Eq. (53) and $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ in Eq. (57) are the same, explaining why these maximum rates agree in scenarios S2-S3 as shown in Figs. 5(a), 6(a), and 7(a). Additionally, for $\kappa=0.1$, Eq. (56) is the dashed line that converges to the $|\sin\theta_{1}|$ value when maximizing $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ in the multi-TeV region.
@@ V Conclusion
We found the maximum resonant production rates of $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}/h_{1}h_{3}/h_{3}h_{3}$ for scenarios including (S1) current constraints on this model and projected constraints from the (S2) HL-LHC, (S3) FCC-ee+HL-LHC, and (S4) ILC500+HL-LHC. In all scenarios, for $h_{2}$ masses below 1 TeV, the various maximum rates are comparable with $h_{3}h_{3}$ having the largest maximum rate followed by $h_{1}h_{3}$ then $h_{1}h_{1}$. As $m_{2}$ increases into the multi-TeV range, some surprising behavior occurs. First, the rate for $pp\rightarrow h_{2}\rightarrow h_{1}h_{3}$ and the corresponding branching ratio $h_{2}\rightarrow h_{1}h_{3}$ decrease quickly becoming quite small at 10 TeV. The maximum rate for $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ and $pp\rightarrow h_{2}\rightarrow h_{3}h_{3}$ converge to the same value, with the corresponding branching ratios ${\rm BR}(h_{2}\rightarrow h_{3}h_{3})\approx 0.5$ and ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})\approx 0.25$. We showed that this behavior can be understood analytically through a combination constraints from requirements of the global minimum, perturbative unitary, narrow width, and Goldstone boson equivalence theorem. Indeed, analytical formula were developed to understand the generality of these behaviors. All benchmark model points corresponding to these maximum rates have been uploaded with the arXiv version of this paper.
@@
Signals with three and four SM-like Higgs bosons in the final state are possible when $m_{3}=270$ GeV. For the parameter points that maximized the $h_{1}h_{3}$ or $h_{3}h_{3}$ rates we found that ${\rm BR}(h_{3}\rightarrow h_{1}h_{1})\approx 1$ for the large majority of points. The three Higgs signal originate from $h_{1}h_{3}$, whose maximum production rate decreases sharply in the multi-TeV range. Since the maximum rates for $h_{3}h_{3}$ and $h_{1}h_{1}$ are similar, we may expect to see both Higgs pair and quartet production when kinematically allowed. In the multi-TeV range, this model has the surprising feature that it could be discovered in the four Higgs and di-Higgs signals before three Higgs.

### v2

@@ IV Benchmark Points
  • S1: We apply the current limits, described in Section III, on $\sin\theta_{1}$ to produce expected maximum rates for the HL-LHC. The most constraining limits on $\sin\theta_{1}$ from Fig. 1 are used. Results from both the hard cut method for scalar searches with Higgs precision data constraints and $\chi^{2}_{\rm Tot}$ fits from Eq. (41) are considered separately. In practice, the strongest constraints on $\sin\theta_{1}$ are found by considering the smallest maximum value of $|\sin\theta_{1}|$ from the ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=0$ and ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})=1$ cases.
@@
Scenario S1 is intended to find maximum di-scalar resonance rates for the HL-LHC. Hence, the $h_{2}$ mass range $m_{i}+m_{j}\leq m_{2}\leq 1$ TeV is used for S1, where $m_{i}$ and $m_{j}$ are the masses of the $h_{2}$ decay products. Scenarios S2-S4 are used to project maximum rates for future $pp$ machines, such as the 100 TeV FCC-hh. In these scenarios we consider a mass range of $m_{i}+m_{j}\leq m_{2}\leq 12$ TeV. For scenarios S2-S4, we normalize to the maximum rates to the SM Higgs production rate at the mass of $h_{2}$. Hence, they are very generic and can be applied to many future collider proposals and many single $h_{2}$ production modes.
@@
While a range of $m_{2}$ are considered, only certain values of the mass of $h_{3}$ are used. These masses are chosen to open new decay channels for $h_{3}$ and possibly new phenomenology for the resonance searches. Assuming $\theta_{2}$ is negligible but non-zero, $h_{3}$ will have partial widths into SM gauge bosons and fermions:
@@
Whether or not these are the dominant branching ratios depends on the Higgs trilinear coupling $h_{1}-h_{1}-h_{3}$ which in turn depends on the parameter point that maximizes the $h_{2}\rightarrow h_{1}h_{3}$ or $h_{2}\rightarrow h_{3}h_{3}$ production rates. As we show below, for all benchmark points, the branching ratio of $h_{3}\rightarrow h_{1}h_{1}$ is at or near one when kinematically allowed.
For the benchmark masses $m_{3}=130$ and $200$ GeV, there are possible decays of $h_{3}$ into off-shell $h_{1}$: $h_{3}\rightarrow h_{1}^{*}h_{1}^{*}$ and $h_{3}\rightarrow h_{1}h_{1}^{*}$. In principle, these decays could compete with the on-shell decays $h_{3}\rightarrow bb$ and $h_{3}\rightarrow WW$. The branching ratios of $h_{3}$ into off-shell $h_{1}$ will depend on the relevant trilinear: $\lambda_{113}$. Upper limits on $\lambda_{113}$ can be found using the bounds on scalar potential parameters in App. B. As discussed below, there is a subtlety on whether or not the rate of $h_{2}\rightarrow h_{1}h_{3}$ or $h_{2}\rightarrow h_{3}h_{3}$ is being maximized. As seen in Eqs. (18,21), in the small angle limit, the leading behavior of $\lambda_{113}$ is the same as $\lambda_{123}$. Hence, if $h_{2}\rightarrow h_{1}h_{3}$ is being maximized, the relevant upper bound on the $h_{1}-h_{1}-h_{3}$ coupling is
$|\lambda_{113}|$ $\lesssim$ $4\sqrt{\frac{2\pi}{3}}m_{1}|\sin\theta_{1}|\left|\cos\theta_{1}\sqrt{1+\frac{m_{2}^{2}-m_{1}^{2}}{m_{1}^{2}}\sin^{2}\theta_{1}}\right.$
$\left.+\sqrt{2}\sin\theta_{1}\frac{m_{2}}{m_{1}}\left(\sqrt{1+\frac{m^{2}_{3}}{m^{2}_{2}}-\frac{m_{2}^{2}-m_{1}^{2}}{m_{2}^{2}}\sin^{2}\theta_{1}}+\frac{m_{3}}{2\,m_{2}}\right)\right|+\mathcal{O}(\sin\theta_{2})$
The rate for $h_{2}\rightarrow h_{3}h_{3}$ is maximized when the rate for $h_{2}\rightarrow h_{1}h_{3}$ is minimized, i.e. $\lambda_{123}=0$. This corresponds to the first term in Eq. (21) being zero. In this case the relevant upper bound on $\lambda_{113}$
$|\lambda_{113}|$ $\lesssim$ $|\sin\theta_{1}\sin\theta_{2}|\left|\frac{2\,m_{1}^{2}+m_{3}^{2}}{v_{\rm EW}}\cos^{2}\theta_{1}+4\sqrt{\frac{2\pi}{3}}m_{1}\left[\left(1-\frac{3}{2}\sin^{2}\theta_{1}\right)\sqrt{1+\frac{m_{2}^{2}-m_{1}^{2}}{m_{1}^{2}}\sin^{2}\theta_{1}}\right.\right.$ (47)
$\left.\left.+\frac{3}{2^{3/2}}\sin 2\theta_{1}\frac{m_{2}}{m_{1}}\sqrt{1-\frac{m_{2}^{2}-m_{1}^{2}}{m_{2}^{2}}\sin^{2}\theta_{1}}\right]\right|+\mathcal{O}(\sin^{2}\theta_{2})$
For a benchmark point we use $\sin\theta_{2}=0.01$. For either of the upper bounds in Eqs. (IV,47), we find that at $m_{3}=130$ GeV the branching ratio of $h_{3}\rightarrow h_{1}^{*}h_{1}^{*}$ is at least four orders of magnitude smaller than ${\rm BR}(h_{3}\rightarrow b\bar{b})$. At $m_{3}=200$ GeV, when the bound in Eq. (IV) is saturated the branching ratio of $h_{3}\rightarrow h_{1}h_{1}^{*}$ is a factor of $2-20$ smaller than $h_{3}\rightarrow WW$, while for the bound in Eq. (47) $h_{3}\rightarrow h_{1}h_{1}^{*}$ is at least four orders of magnitude smaller than $h_{3}\rightarrow WW$. Hence, $h_{3}\rightarrow b\bar{b}$ is the dominate decay at $m_{3}=130$ GeV and $h_{3}\rightarrow WW$ dominates at $m_{3}=200$ GeV. Note that $h_{3}$ couplings to SM fermions and gauge bosons are suppressed by the combination $\sin\theta_{1}\sin\theta_{2}$. The bound in Eq. (IV) is suppressed by only $\sin\theta_{1}$. Hence, whether $h_{3}\rightarrow h_{1}h_{1}^{*}$ or $h_{3}\rightarrow WW$ are dominant at $m_{3}=200$ GeV can be sensitive to the choice of $\sin\theta_{2}$ when $h_{2}\rightarrow h_{1}h_{3}$ is being maximized. At $m_{3}=130$ GeV, $h_{3}\rightarrow bb$ is expected to always be dominant due to the very large phase space and $h_{1}$ coupling constant suppression in $h_{3}\rightarrow h_{1}^{*}h_{1}^{*}$. Finally, the bound in Eq. (47) is suppress by $\sin\theta_{1}\sin\theta_{2}$ and all results derived from that bound is robust against different choices of $\sin\theta_{2}$.
We now discuss the results for maximizing $h_{i}h_{j}$ resonant production under the various scenarios. After these results, we will comment on the decays of $h_{3}$ and the possibility of having three or four SM-like Higgs boson signals.
@@ IV.1.1 Scenario S1: HL-LHC projections
Figure 4: Benchmarks for HL-LHC from scenario S1, showing (a,c) maximum production rates for $pp\rightarrow h_{2}\rightarrow h_{i}h_{j}$ with the gluon fusion and VBF modes added together at $\sqrt{S}=14$ TeV and (b,d) the corresponding branching ratios to different di-scalar states. Results from using the bounds from $\Delta\chi^{2}_{\rm Tot}$ fits in Eq. (41) are shown in (a,b) and from the hard cut method in (c,d). Blue is for $h_{2}\rightarrow h_{1}h_{1}$, red for $h_{2}\rightarrow h_{1}h_{3}$, and green for $h_{2}\rightarrow h_{3}h_{3}$. Solid is for $m_{3}=130$ GeV, dotted for $m_{3}=200$ GeV, and dashed for $m_{3}=270$ GeV.
In figure 4 we show the upper bounds on (a,c) the maximum production rate for $pp\rightarrow h_{2}\rightarrow h_{i}h_{j}$ and (b,d) the corresponding branching ratios of $h_{2}\rightarrow h_{i}h_{j}$, $i,j=1,3$, for scenario S1. Figure 4(a,b) use constraints from a $\chi^{2}_{\rm Tot}$ fit and (c,d) from the hard cut method for scalar searches with precision Higgs constraints. The maximum production rates and branching ratios are largely independent of the choice of $m_{3}$ once kinematically allowed.
For $m_{2}\lesssim 600$ GeV, the maximum branching ratios for $h_{2}\rightarrow h_{1}h_{3}$ and $h_{2}\rightarrow h_{3}h_{3}$ are near one, with the final state $h_{3}h_{3}$ slightly larger. For $h_{2}\rightarrow h_{1}h_{1}$ the maximum branching ratio is between $0.6$ and $0.8$. As can be seen in Eqs. (II,-19), the values of the $\lambda_{123}$ and $\lambda_{233}$ trilinear couplings have much more freedom from the choices of potential parameters than $\lambda_{112}$. Indeed, the ratio of the $h_{2}\rightarrow h_{1}h_{1}$ and $h_{2}\rightarrow ZZ$ partial widths are
@@
Hence, the partial width of $h_{2}\rightarrow h_{1}h_{1}$ cannot be too far from $h_{2}\rightarrow ZZ$, limiting how large the branching ratio of $h_{2}\rightarrow h_{1}h_{1}$ can be.
Once the mass of $h_{2}$ surpasses 600 GeV, all branching ratios decrease. The partial widths of $h_{2}$ into the massive SM gauge bosons $W^{\pm}W^{\mp}$ and $ZZ$ grow like the cubic power of $m_{2}$, whereas the partial widths into $h_{1}h_{3}$ and $h_{3}h_{3}$ do not. Additionally, the branching ratio of $h_{2}\rightarrow h_{1}h_{1}$ approaches the Goldstone boson equivalence theorem limit of ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})\approx 0.25$. Hence, the maximal branching ratio into $h_{i}h_{j}$ decreases in general. It is worth noting that the maximum branching ratio of the asymmetric decay $h_{2}\rightarrow h_{1}h_{3}$ decreases more quickly than the symmetric decay of $h_{2}\rightarrow h_{3}h_{3}$. In the small angle limit $|\theta_{1}|\ll 1$, the leading term of the trilinear $\lambda_{123}$ depends on trilinear couplings while $\lambda_{233}$ depends on quartic couplings. As we show in Appendix C, the combination of requiring the global minimum be the SM vacuum, perturbative unitarity, and a narrow width places an upper bound on $\lambda_{123}$ that is more stringent than the upper bound on $\lambda_{233}$. Indeed, as shown in Eq. (113), these constraints force the branching ratio of $h_{2}\rightarrow h_{1}h_{3}$ to approach zero as the mass of $h_{2}$ increases. While these derived constraints may not be the most stringent, they do help explain qualitatively why the branching ratio of $h_{2}\rightarrow h_{1}h_{3}$ decreases more quickly than the other di-scalar modes as $m_{2}$ increases.
@@ IV.1.2 Scenarios S2-S4: Future Collider Projections
The fermionic partial widths only grow linearly with $m_{2}$ while the bosonic widths grow as $m_{2}^{3}$. Hence, $h_{2}\rightarrow f\bar{f}$, where $f$ are SM fermions, can be neglected and ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})\approx 0.25$.
Refer to captionRefer to caption
@@
Two particularly surprising things occur when $m_{2}$ is in the multi-Tev range and above: the branching ratio of $h_{2}\rightarrow h_{3}h_{3}$ at its maximum rate approaches 0.5 and the maximum rates of $h_{2}\rightarrow h_{1}h_{1}$ and $h_{2}\rightarrow h_{3}h_{3}$ approach the same value. To understand ${\rm BR}(h_{2}\rightarrow h_{3}h_{3})\approx 0.5$, note that we are maximizing the rate
@@
where we have used the narrow width approximation. Assuming the rate $pp\rightarrow h_{2}\rightarrow h_{3}h_{3}$ is maximized then ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})=0$ and the fermionic branching ratios are negligible. Equation (50) can be rewritten as
@@
A similar calculation can be made for $pp\rightarrow h_{1}h_{3}$, which explains why $|\sin\theta_{1}|$ at the maximum rate agrees for $h_{1}h_{3}$ and $h_{3}h_{3}$ production [see Fig. 3]. As discussed previously, theoretical constraints force ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})\rightarrow 0$ as $m_{2}$ becomes large. Hence, branching ratios and rates in Eqs. (54,55) cannot be saturated. There are constraints on $\lambda_{233}$ from global minimum, perturbative unitarity, and narrow width as shown in Eq. (103). However, these are significantly weaker than for $\lambda_{123}$, and the $pp\rightarrow h_{2}\rightarrow h_{3}h_{3}$ can saturate the bounds in Eqs. (54,55).
Figure 6: Benchmarks for future colliders, constrained by projections for HL-LHC and FCC-ee in scenario S3, showing (a) the maximum production times branching ratio scaled to the SM-like Higgs production rate and (b) the corresponding branching ratios. Blue is for $h_{2}\rightarrow h_{1}h_{1}$, red for $h_{2}\rightarrow h_{1}h_{3}$, and green for $h_{2}\rightarrow h_{3}h_{3}$. Solid is for $m_{3}=130$ GeV, dotted for $m_{3}=200$ GeV, and dashed for $m_{3}=270$ GeV.
@@
The maximum rates for $pp\rightarrow h_{2}\rightarrow h_{3}h_{3}$ in Eq. (55) and $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ in Eq. (59) are the same, explaining why these maximum rates agree in scenarios S2-S4 as shown in Figs. 5(a), 6(a), and 7(a). Additionally, for $\kappa=0.1$, Eq. (58) is the dashed line that converges to the $|\sin\theta_{1}|$ value when maximizing $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ in the multi-TeV region.
@@ V Conclusion
We found the maximum resonant production rates of $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}/h_{1}h_{3}/h_{3}h_{3}$ for scenarios including (S1) current constraints on this model and projected constraints from the (S2) HL-LHC, (S3) FCC-ee+HL-LHC, and (S4) ILC500+HL-LHC. In all scenarios, for $h_{2}$ masses below 1 TeV, the various maximum rates are comparable with $h_{3}h_{3}$ having the largest maximum rate followed by $h_{1}h_{3}$ then $h_{1}h_{1}$. As $m_{2}$ increases into the multi-TeV range, some surprising behavior occurs. First, the rate for $pp\rightarrow h_{2}\rightarrow h_{1}h_{3}$ and the corresponding branching ratio $h_{2}\rightarrow h_{1}h_{3}$ decrease quickly becoming quite small at 10 TeV. The maximum rate for $pp\rightarrow h_{2}\rightarrow h_{1}h_{1}$ and $pp\rightarrow h_{2}\rightarrow h_{3}h_{3}$ converge to the same value, with the corresponding branching ratios ${\rm BR}(h_{2}\rightarrow h_{3}h_{3})\approx 0.5$ and ${\rm BR}(h_{2}\rightarrow h_{1}h_{1})\approx 0.25$. We showed that this behavior can be understood analytically through a combination of constraints from requirements of the global minimum, perturbative unitary, narrow width, and Goldstone boson equivalence theorem. Indeed, analytical formulas were developed to understand the generality of these behaviors. All benchmark model points corresponding to these maximum rates have been uploaded with the arXiv version of this paper.
@@
Signals with three and four SM-like Higgs bosons in the final state are possible when $m_{3}=270$ GeV. For the parameter points that maximized the $h_{1}h_{3}$ or $h_{3}h_{3}$ rates we found that ${\rm BR}(h_{3}\rightarrow h_{1}h_{1})\approx 1$ for the large majority of points. The three Higgs signal originates from $h_{1}h_{3}$, whose maximum production rate decreases sharply in the multi-TeV range. Since the maximum rates for $h_{3}h_{3}$ and $h_{1}h_{1}$ are similar, we may expect to see both Higgs pair and quartet production when kinematically allowed. In the multi-TeV range, this model has the surprising feature that it could be discovered in the four Higgs and di-Higgs signals before three Higgs.